OpenStudy (anonymous):

If P(x) = ax4+bx3+cx2+dx+e has roots at x = 1, 2, 3, 4 and P(0) = 48, what is P(5)

OpenStudy (anonymous):

48

OpenStudy (anonymous):

now you have p(0)=48 p(1),p(2),p(3),and p(4)=0 there are five constants (unknowns ) and five equations solve these equations together to get the unknowns

OpenStudy (anonymous):

p(5)=48

OpenStudy (anonymous):

cuase p(0) =48 so we have the four roots (x-1) * (x-2) *(x-3) *(x-4) the four root so p(0)= -1*-2*-3*-4 = 24 so we need 24 that p(x) =(x-1) * (x-2) *(x-3) *(x-4) +24 p(5)= (5-1) * (5-2) *(5-3) *(5-4)+24 = 48